The equation of the line L is 3y - x=9. Find
a) The coordinates of the point where l intersects the y- axis.
b) The equation of the line which is parallel to l and which passes through the point (0, -5)

Answer by Alan3354(21583)

The equation of the line L is 3y - x=9. Find
a) The coordinates of the point where l intersects the y- axis.
Set x = 0 to get the y-intercept, and vice versa.
3y = 9
y = 3
y-intercept = (0,3)
-------------
-x = 9
x-intercept = (-9,0)
------------------------
b) The equation of the line which is parallel to l and which passes through the point (0,-5)
Find the slope, m, of the line by putting it into slope-intercept form. That means solve for y.
3y-x = 9
3y = x+9
y = (1/3)x + 3
m = 1/3
Lines parallel have the same slope, m = 1/3
Use
y-y1 = m*(x-x1) where (x1,y1) is the point.
y+5 = (1/3)x
y = (1/3)x - 5 slope-intercept form
x - 3y = 15 standard form